# arXiv:2603.00624v1[cs.LG] 28 Feb 2026

### v1

@@ 3.3 Standard Idempotent Module:Training the Network Idempotent
where $\mathcal{T}_{t}$ is the current task and $y^{*}$ is the second input that is set to either the ground-truth one-hot vector $y$ or the neutral "empty" signal input $\mathbf{0}$. We randomly select $y$ with probability 1P1-P and the neutral "empty" signal input $\mathbf{0}$ with probability $P$. The empty signal $\mathbf{0}$ is defined as a uniform distribution over all classes.
@@ 4.2 Additional Analysis
Cross-Platform Performance. To assess the hardware sensitivity of our results, we run ER+ID with the same configuration on two different platforms: NVIDIA RTX 4090 and Ascend 910B. As Table 6 shows, the performance is consistent across platforms: the differences are small suggesting that the gains from IDER are not hardware-specific. Minor discrepancies are expected due to differences in low-level kernels, numerical precision, and non-determinism in deep learning libraries.
@@
Table 6: Performance comparison for ER+ID between ASCEND and NVIDIA. All experiments are repeated 5 times with different seeds.
@@ ACKNOWLEDGEMENT
This project is supported by the National Natural Science Foundation of China (No. 62406192), Shanghai Municipal Science and Technology Major Project (No. 2025SHZDZX025G03), Opening Project of the State Key Laboratory of General Artificial Intelligence (No. SKLAGI2024OP12), the Tencent WeChat Rhino-Bird Focused Research Program, Kuaishou Technology, and the SJTU Kunpeng & Ascend Center of Excellence.
@@ D.4 Complexity and Training Cost
Appendix E Statement on the Use of Large Language Models
Appendix F Limitations
@@
Appendix G Ablation Results
G.1 contribution of each component
@@ G.2 Hyperparameter sensitivity
G.3 Probability selection
We provide ablation study of the probability $P$ used in the Standard Idempotent Module, which determines whether the second input is set to the empty signal or the ground-truth one-hot vector. We perform a sensitivity ablation on Split CIFAR-100 with 500 buffer size and on ER+ID method. In our ablation study, the second input is set to the neutral "empty" signal input 0 with probability $P$ and the ground-truth one-hot vector y with probability 1P1-P. Table 10 shows that as $P$ increases, FAA consistently improves and peaks at $P=0.9$, which reaches 44.82. Then there is a slight drop at $P=1.0$, where FAA is 43.26. Therefore, we use $P=0.9$ by default. For simplicity, we can also set $P=1.0$, which yields comparable performance.

### v2

# arXiv:2603.00624v2[cs.LG] 03 Mar 2026

@@ 3.3 Standard Idempotent Module:Training the Network Idempotent
where $\mathcal{T}_{t}$ is the current task and $y^{*}$ is the second input that is set to either the ground-truth one-hot vector $y$ or the neutral “empty” signal input $\mathbf{0}$. We randomly select $y$ with probability 1P1-P and the neutral “empty” signal input $\mathbf{0}$ with probability $P$. The empty signal $\mathbf{0}$ is defined as a uniform distribution over all classes.
@@ 4.2 Additional Analysis
Cross-Platform Performance. To assess the hardware sensitivity of our results, we run ER+ID with the same configuration on two different platforms: NVIDIA RTX 4090 and Huawei Ascend 910B. As Table 6 shows, the performance is consistent across platforms: the differences are small suggesting that the gains from IDER are not hardware-specific. Minor discrepancies are expected due to differences in kernels, numerical precision, and non-determinism in deep learning libraries.
@@
Table 6: Performance comparison for ER+ID between NVIDIA RTX 4090 and Huawei Ascend 910B. All experiments are repeated 5 times with different seeds.
@@ ACKNOWLEDGEMENT
This project is supported by the National Natural Science Foundation of China (No. 62406192), the Shanghai Municipal Special Program for Basic Research on General AI Foundation Models (Grant No. 2025SHZDZX025G03), Opening Project of the State Key Laboratory of General Artificial Intelligence (No. SKLAGI2024OP12), the Tencent WeChat Rhino-Bird Focused Research Program, Kuaishou Technology, and the SJTU Kunpeng & Ascend Center of Excellence.
@@ D.4 Complexity and Training Cost
Appendix E Limitations
@@
Appendix F Ablation Results
F.1 contribution of each component
@@ F.2 Hyperparameter sensitivity
F.3 Probability selection
We provide ablation study of the probability $P$ used in the Standard Idempotent Module, which determines whether the second input is set to the empty signal or the ground-truth one-hot vector. We perform a sensitivity ablation on Split CIFAR-100 with 500 buffer size and on ER+ID method. In our ablation study, the second input is set to the neutral “empty” signal input 0 with probability $P$ and the ground-truth one-hot vector y with probability 1P1-P. Table 10 shows that as $P$ increases, FAA consistently improves and peaks at $P=0.9$, which reaches 44.82. Then there is a slight drop at $P=1.0$, where FAA is 43.26. Therefore, we use $P=0.9$ by default. For simplicity, we can also set $P=1.0$, which yields comparable performance.